You earn $9 per hour. On major holidays, such as Thanksgiving, you earn twice as much per hour. You earned a total of $405 for the week including thanksgiving. Write and solve an equation to find how many hours you worked on thanksgiving if you worked 35 hours during the rest of the week.
Answer by mananth(16949)

thanks giving day x hours
..
other days 35-x hours
..
earning = 9*(35-x) on normal day
Thanksgiving day - 18x
..
18x+9(35-x)=405
18x+315-9x = 405
9x=405-315
9x=90
x=10 hours on Thanksgiving day